OpenStudy (anonymous):

Help for MEDAL : question about sequences n 1 2 3 4 5 6 f(n) 1 2 3 5 8 13

OpenStudy (anonymous):

Which recursive function best represents the values shown in the table? f(1) = 1, f(2) = 2, f(n) = f(n -1) f(n - 2); n > 2 f(1) = 1, f(2) = 2, f(n) = f(n -1) + f(n - 2); n > 2 f(1) = 1, f(2) = 2, f(n) = f(n -3) + f(n - 2); n > 2 f(1) = 1, f(2) = 2, f(n) = f(n -3) f(n - 2); n > 2

OpenStudy (anonymous):

I mean i dont even think its either a geometric sequence or arithmetic

OpenStudy (anonymous):

ill help you dont worry

OpenStudy (anonymous):

If you need the explicit formulas let me know

OpenStudy (anonymous):

yes i would like that it would be very help ful

OpenStudy (anonymous):

Arithmetic explicit formula: f(n)=f(1)+d(n-1) Geometric explicit formula : f(n)=f(1)*r ^n-1

OpenStudy (anonymous):

\[f(n)=f(1)*r ^{n-1}\]

OpenStudy (anonymous):

thats the geometric , just to make it easier to see

OpenStudy (anonymous):

hold on im going to get some help

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

Hey sorry was afk but what guy commented? he must of deleted the comments. Thanks everyone who tried to help :)

OpenStudy (gorv):

f(1)=1 f(2)=2

OpenStudy (gorv):

f(n)=f(n-1)+f(n-2) for n>2 as for n=3 f(3)=f(3-1)+f(3-2)=f(2)+f(1)=2+1=3

OpenStudy (gorv):

2nd one

OpenStudy (anonymous):

Thank you! @gorv